# Can We Hide Machines in the Crowd? Quantifying Equivalence in LLM-in-the-loop Annotation Tasks

### v1

@@ Abstract.
Many evaluations of large language models (LLMs) in text annotation focus primarily on the correctness of the output, typically comparing model-generated labels to human-annotated “ground truth” using standard performance metrics. In contrast, our study moves beyond effectiveness alone. We aim to explore how labeling decisions–by both humans and LLMs–can be statistically evaluated across individuals. Rather than treating LLMs purely as annotation systems, we approach LLMs as an alternative annotation mechanism that may be capable of mimicking the subjective judgments made by humans. To assess this, we develop a statistical evaluation method based on Krippendorff’s $\alpha$, paired bootstrapping, and the Two One-Sided t-Tests (TOST) equivalence test procedure. This evaluation method tests whether an LLM can blend into a group of human annotators without being distinguishable.
We apply this approach to two datasets–MovieLens 100K and PolitiFact–and find that the LLM is statistically indistinguishable from a human annotator in the former ($p=0.004$), but not in the latter ($p=0.155$), highlighting task-dependent differences. It also enables early evaluation on a small sample of human data to inform whether LLMs are suitable for large-scale annotation in a given application.
@@ 1. Introduction
In his landmark 1950 paper, Alan Turing proposed a criterion for machine intelligence: if a machine can engage in conversation such that a human evaluator cannot reliably distinguish it from another human. The machine can be said to exhibit intelligent behaviour (Turing, 2009). This formulation–now known as the Turing Test–shifts the focus from how a machine works inside to how it behaves on the outside.
@@
In this work, we build on the spirit of the Turing Test. However, we focus on a more domain-specific setting: text annotation tasks. Instead of testing a machine through open-ended conversations, we ask a simpler question. Can an LLM act like a human annotator? Specifically, can its output be statistically indistinguishable from that of people in a multi-person annotation task?
@@
While LLMs have shown strong performance on many general-purpose classification tasks (Nasution and Onan, 2024), their role in subjective or domain-specific annotation settings remains unclear. In information retrieval contexts—such as document relevance assessment, intent classification, or stance detection—annotation often involves subtle, context-dependent judgments (Faggioli et al., 2023; Roitero et al., 2023; Kutlu et al., 2018a). These applications continue to rely on traditional classifiers trained on human-labeled data, particularly where interpretability, auditability, or fairness are required (Lin et al., 2021; MacAvaney and Soldaini, 2023). If LLMs can produce labels that are indistinguishable from those generated by humans in such scenarios, they offer a path to reducing annotation costs while preserving human-like interpretive behavior. This makes it important to understand whether LLMs can label correctly and effectively participate in the human annotation process used to train these models.
@@
Text classification is widely used for natural language processing (NLP) tasks, such as news categorization, sentiment analysis, and subject labeling (Dong et al., 2022; Siino et al., 2025). It involves assigning labels to textual elements such as sentences, questions, paragraphs, or entire documents. Some of the classifications require human labeling to train and test the machine learning models. Annotations act as the ground truth against which models are tested, refined, and advanced.
@@
Annotation tasks often depend on subjective human judgment rather than a single, objective truth. For example, in relevance assessments, people may interpret the same content differently, and their judgments can change over time (Faggioli et al., 2023). So, the aim is not always to identify the “correct” label, but to understand how labeling decisions emerge across different individuals. LLMs are increasingly good at producing convincing outputs. However, it is still unclear whether their responses reflect human judgment, especially in the absence of evidence that their outputs are rooted in actual human experience (Faggioli et al., 2023).
@@
Traditional human annotation can be constrained by cost and consistency (Devillers et al., 2005). Recent research has compared the quality of annotations by LLM and human in use of NLP application (Nasution and Onan, 2024). Also, study has investigated using LLMs to help or even replace human annotators on some tasks (Alaofi et al., 2024), usually by comparing their agreement with human results using measures like Krippendorff’s $\alpha$ or Cohen’s kappa. Most of these studies treat the LLM as a single, standalone system and check how well it matches a human-created “ground truth” (MacAvaney and Soldaini, 2023; Alaofi et al., 2024; Zendel et al., 2024; Dietz et al., 2025; Balog et al., 2025).
@@
This work treats LLMs not just as tools for text classification, but as participants that can imitate the subjective—and sometimes inconsistent—judgments made by humans. We introduce a practical method based Krippendorff’s $\alpha$, bootstrapping and TOST to test whether an LLM can blend into a group of human annotators without being identified. This approach only requires a small number of annotation items and functions as a domain-specific take on the Turing Test. It supports early-stage evaluation on a small sample to determine the suitability of LLMs for large-scale annotation. We apply it to a real-world classification task and examine the results.
Our key contributions are:
@@
  • We demonstrate the application of our methodology on two datasets—MovieLens 100K and PolitiFact—showing that the LLM is statistically indistinguishable from human annotators in the former ($p=0.004$) but not in the latter ($p=0.155$), revealing important task-dependent differences.
@@ 2.1. Text Classification
Text classification is a core task in NLP. It helps organize unstructured text from sources like messages, documents, and websites. To make sense of all this text, researchers and developers use models that automatically classify text into categories like topic, sentiment, or author identity. Siino et al. (2025) provide an overview of common datasets used in these tasks, ranging from author profiling and news categorization to sentiment analysis.
Labeled data is essential for building and testing models, with human annotations serving as the “ground truth” (Nasution and Onan, 2024; Siino et al., 2025). Manual labeling is often costly, slow, and inconsistent, making automation preferable. Models are typically trained on a subset of labeled data and evaluated by comparing their predictions to human annotations.
Recently, LLMs have shown impressive performance across many NLP tasks  (Brown et al., 2020), raising the possibility of replacing traditional supervised models or even human annotators. However, this shift is far from complete (Movva et al., 2024).
Many domain specific real-world applications such as medical coding, legal triage, or sentiment analysis continue to rely on traditional classifiers trained on large datasets labeled by experts. This is because LLMs alone often lack the nuanced understanding required for these tasks. As a result, substantial human annotation remains essential for developing and validating reliable models. In areas where complex human judgment is critical, direct human involvement remains indispensable Christoforou et al. (2025).
@@ 2.2. Human Annotation and Generative AI
Human annotation comes with challenges, most notably cost and consistency (Devillers et al., 2005). Text classification models rely on large amounts of labeled data, and hiring workers to do all the labeling is not always feasible. That is why crowdsourcing has become a solution. Platforms that connect requesters with crowd workers make it possible to outsource labeling tasks like relevance judgments, sentiment tagging, and topic categorization to non-experts (Alonso and Mizzaro, 2012; Kutlu et al., 2018a). This approach has been crucial in building datasets needed to train machine learning models.
Crowdsourcing was once seen as flexible and empowering for workers. However, it is often criticized as invisible, low-paid labor that supports modern AI behind the scenes.
Given these concerns, researchers have started asking: Can generative AI (GenAI), especially LLMs, step in and take over some of these annotation tasks? Some early findings suggest that LLMs tend to do well on straightforward tasks like summarization or basic sentiment analysis (Christoforou et al., 2025). But when the task requires more nuanced judgment—like interpreting sarcasm, ambiguity, or subtle context—human annotators still outperform the machines (Nasution and Onan, 2024).
So, the real question is not just whether GenAI can get the “righ” answer. It is whether its decisions reflect the kinds of judgments humans would make, especially in cases where there is no single correct label. One study (Faggioli et al., 2023) looked at how well LLMs could handle relevance judgments, a task where subjectivity plays a big role. While LLMs showed some ability to mimic human responses, they were not consistent or nuanced enough to fully replace human workers.
@@ 2.3. Existing Evaluation Method
This raises an important question: if an LLM’s annotation looks like a human’s, does that mean it is truly the same—or are we missing deeper differences in how judgments are made? In many tasks, there is no single “correct” answer; human judgments are often subjective, context-dependent, and inconsistent over time (Bailey et al., 2008; Faggioli et al., 2023).
To better reflect this, we propose a new approach: instead of checking if an LLM agrees with the crowd, we ask whether it can blend into the crowd—becoming statistically indistinguishable from human annotators, and that we call it a “Statistical Turing Test”.
@@ 2.4. Human Judgment
This inconsistency can be attributed to the variability in cue weights applied across different tasks. Prior research has shown that judgment consistency tends to decrease as the number of cues increases (Einhorn, 1971). In contrast, LLMs often display greater consistency in annotation tasks (Faggioli et al., 2023), likely due to more stable internal representations of cues and weights.
To evaluate whether the LLM’s cue integration falls within an acceptable threshold of variability, we use inter-annotator agreement (IAA) as a proxy for measuring consistency in annotation judgments.
@@ 2.5. Inter-Annotator Agreement (IAA)
Researchers who rely on hand-labeled data–where items are manually labeled with categories for empirical analysis or model development–must demonstrate that the labeling process is reliable (Artstein and Poesio, 2008). A fundamental assumption in annotation methodology is that the data are considered reliable when multiple annotators agree on the labels assigned, to a degree appropriate for the objectives of the study (Krippendorff, 2018; Craggs and Wood, 2004). Consistent agreement among annotators suggests that they share a common understanding of the annotation guidelines, and thus can be expected to apply those guidelines consistently.
IAA is a standard metric used to quantify this consistency. In multi-people annotation settings, annotators may have varied backgrounds and limited domain expertise. IAA helps determine whether labels are trustworthy and whether a task is clearly defined or inherently subjective. High agreement indicates clear instructions and straightforward data, while low agreement may reveal task ambiguity, multiple valid interpretations, or inconsistent annotator behavior.
@@ 3. Methodology
Building on the idea discussed in Section 2.5, consistent labeling by multiple human annotators suggests that they share a common understanding of the annotation guidelines and apply them reliably (Artstein and Poesio, 2008). Inspired by the logic of the Turing Test (Turing, 2009), we propose a methodology that uses inter-annotator agreement to evaluate whether a large language model (LLM) can function as an individual annotator—that is, whether it can serve as a substitute for a human in the annotation process.
@@ 3.1.1. Motivation: Substituting Human Annotators
Krippendorff’s $\alpha$ measures the extent of agreement among annotators using the formula: $\alpha=1-\frac{D_{o}}{D_{e}}$, where $D_{o}$ is the observed disagreement and $D_{e}$ is the expected disagreement by chance.
@@ 3.1.3. Measuring Agreement Across Substitutions
We begin by computing Krippendorff’s $\alpha$ for the original group of $i$ human annotators. We denote this baseline agreement as $\alpha_{1}=\text{Krippendorff's $\alpha$ for human group}$. Next, we compute $\alpha$ for each of the $i$ modified groups where one human annotator is replaced by the LLM. These are denoted as $\alpha_{j}=\text{Krippendorff’s $\alpha$ for substitution group}j=2,\dots,i+1$.
@@ 3.1.4. Paired Bootstrap for Variability Estimation
Let $\alpha_{1}=\left\{\alpha_{1}^{(1)},\alpha_{1}^{(2)},\dots,\alpha_{1}^{(B)}\right\}$ represent the $\alpha$ values computed for the original human group across the $B$ bootstrap samples. For each LLM-substituted group $j=2,\dots,i+1$, we similarly compute a distribution of alpha values: $\alpha_{j}=\left\{\alpha_{j}^{(1)},\alpha_{j}^{(2)},\dots,\alpha_{j}^{(B)}\right\}$.
@@ 3.2. Equivalence Testing with TOST
The null hypothesis is that the difference exceeds the equivalence margin: $H_{0}:|x_{1}-x_{2}|>\Delta_{\text{equiv}}$. We reject $H_{0}$ if both $t_{1}$ and $t_{2}$ fall within the critical region for their respective one-sided tests, thereby concluding that the observed difference is within an acceptable range of equivalence.
@@ 3.2.1. Equivalence Margin Definition
The equivalence margin $\Delta_{\text{equiv}}$ is a threshold below which the difference is considered negligible. We estimate this margin empirically based on natural variability among human annotators: $\Delta_{\text{equiv}}=(\alpha_{a}-\alpha_{b})\cdot\text{fraction}$, where:
@@ 3.3.2. Annotator Group Size
We analyze how Krippendorff’s $\alpha$ changes when one human annotator is replaced by an LLM. Let $i$ be the number of annotators, $n$ the number of items, and $\ell_{ak}$ the label from annotator $a$ on item $k$. As mentioned in Section 2.6, Krippendorff’s $\alpha$ is defined as $\alpha=1-\frac{D_{o}}{D_{e}}$, where $D_{o}$ is the average observed pairwise disagreement and $D_{e}$ is the expected disagreement under random labeling based on marginal label distributions.
@@ 4.1. Dataset
MovieLens 100K (Harper and Konstan, 2015) Movie ratings 1 682
WebCrowd25k (Kutlu et al., 2018b) IR relevance $\approx$4 500
TREC-8 Re-assessments (Roitero et al., 2021) IR relevance 4 269
@@
MBIC (Spinde et al., 2021) Media bias $\approx$2 600
CoQA (Reddy et al., 2019) Conversational QA 8 000
POPQUORN (Pei and Jurgens, 2023) Offensive–QA $\approx$5 500
D3CODE (Mostafazadeh Davani et al., 2024) Cross-cultural offensiveness detection $\approx$4 500
@@
CODA-19 (Huang et al., 2020) COVID-19 abstract section labeling 10 966
We compiled a list of publicly available datasets that provide multiple annotations per item along with identifiable annotator IDs, enabling us to trace which individual labeled each item and to perform annotator-level substitutions. These datasets are summarized in Table 1.
To fit our evaluation experiments, we filtered the datasets using a set of selection criteria defined by our evaluation methodology. The code for this filtering process is available. [^2] Our filtering code can be found in https://github.com/peanutH/LLM-evaluation. Specifically, we required that the dataset contain two disjoint groups of annotators, each of size $i$, such that there are at least 80 items, with each item annotated by at least two annotators from each group. After data filtering, two datasets satisfy our criteria: MovieLens 100K (Harper and Konstan, 2015) and PolitiFact (Roitero et al., 2020).
@@
The original PolitiFact dataset collected by Wang (2017) contains 12,000 statements produced by U.S. politicians, each statement is labeled by an expert judge on a six-level scale for the statement’s truthfulness. Roitero et al. (2020) selected 120 statements, 20 for each truth level, related to COVID-19 from the original PolitiFact dataset. Then, workers are recruited from the Amazon Mechanical Turk platform to annotate each statement. Overall, each statement was annotated by 10 workers over three different scales: three-level, six-level, one-hundred level. In this work, we use the three-level dataset.
To obtain the LLM annotations, we used GPT-4o mini to annotate the same set of items selected for this experiment, as described in Section 4.2. Below is our LLM prompt for generating annotations for the MovieLens dataset [^3] Our LLM prompts and generated annotations available at https://github.com/peanutH/LLM-evaluation:
--SYSTEM--
You are an average movie watcher. Rate each movie
from 1 to 5 based on how much you liked it overall.
Consider the story, acting, and overall enjoyment.
--USER--
Respond with a list of ratings, one for each movie,
in the same order as presented. Only include the
numeric ratings and nothing else.
@@
{list_of_movies}
The prompts were designed to instruct the LLM to follow the same annotation guidelines provided to human annotators.

### v2

@@ Abstract.
Many evaluations of large language models (LLMs) in text annotation focus primarily on the correctness of the output, typically comparing model-generated labels to human-annotated “ground truth” using standard performance metrics. In contrast, our study moves beyond effectiveness alone. We aim to explore how labeling decisions–by both humans and LLMs–can be statistically evaluated across individuals. Rather than treating LLMs purely as annotation systems, we approach LLMs as an alternative annotation mechanism that may be capable of mimicking the subjective judgments made by humans. To assess this, we develop a statistical evaluation method based on Krippendorff’s $\alpha$, paired bootstrapping, and the Two One-Sided t-Tests (TOST) equivalence test procedure. This evaluation method tests whether an LLM can blend into a group of human annotators without being distinguishable. We apply this approach to two datasets, MovieLens 100K and PolitiFact, and find that the LLM is statistically indistinguishable from a human annotator in MovieLens 100K ($p=0.004$), but not in PolitiFact ($p=0.155$), highlighting task-dependent differences.
@@ 1. Introduction
In Alan Turing’s 1950 landmark 1950 paper, he proposed a criterion for machine intelligence: if a machine can engage in conversation such that a human evaluator cannot reliably distinguish it from another human. The machine can be said to exhibit intelligent behaviour (Turing, 2009). This formulation, now known as the Turing Test, shifts the focus from a machine’s internal workings to its external behaviour. In this work, we build on the spirit of the Turing Test. However, we focus on a more domain-specific setting: text annotation tasks. Instead of testing a machine through open-ended conversations, we ask a simpler question. Can an LLM act like a human annotator? Specifically, can its output be statistically indistinguishable from that of people in a multi-person annotation task?
@@
While LLMs have shown strong performance on many general-purpose classification tasks (Nasution and Onan, 2024; Zhang et al., 2024; Hegselmann et al., 2023), their role in subjective or domain-specific annotation settings remains unclear. In information retrieval contexts—such as document relevance assessment, intent classification, or stance detection—annotation often involves subtle, context-dependent judgments (Faggioli et al., 2023; Roitero et al., 2023; Kutlu et al., 2018a; Schnabel et al., 2025). These applications continue to rely on traditional classifiers trained on human-labeled data, particularly where interpretability, auditability, or fairness are required (Lin et al., 2022; MacAvaney and Soldaini, 2023). If LLMs can produce labels that are indistinguishable from those generated by humans in such scenarios, they offer a path to reducing annotation costs while preserving human-like interpretive behavior. This makes it crucial to understand whether LLMs can accurately and effectively participate in the human annotation process used to train these models.
@@
Text classification is widely used for natural language processing (NLP) tasks, such as news categorization, sentiment analysis, or subject labeling (Dong et al., 2024; Siino et al., 2025). It involves assigning labels to textual elements such as sentences, questions, paragraphs, or entire documents. Some of the classifications require human labeling to train and test the machine learning models. Annotations act as the ground truth against which models are tested, refined, and advanced.
@@
Annotation tasks often depend on subjective human judgment rather than a single, objective truth (Magnossão de Paula et al., 2025). For example, in relevance assessments, people may interpret the same content differently, and their judgments can change over time (Faggioli et al., 2023). So, the aim is not always to identify the “correct” label, but to understand how labeling decisions emerge across different individuals. LLMs are increasingly good at producing convincing outputs. However, it is still unclear whether their responses reflect human judgment, especially in the absence of evidence that their outputs are rooted in actual human experience (Faggioli et al., 2023).
@@
Traditional human annotation can be constrained by cost and consistency (Devillers et al., 2005). Recent research has compared the quality of annotations by LLM and humans for NLP applications (Nasution and Onan, 2024). Also, a study investigated using LLMs to help or even replace human annotators on some tasks (Alaofi et al., 2024), usually by comparing their agreement with human results using measures like Krippendorff’s $\alpha$ or Cohen’s kappa. Most of these studies treat the LLM as a single, standalone system and check how well it matches a human-created “ground truth” (MacAvaney and Soldaini, 2023; Alaofi et al., 2024; Zendel et al., 2024; Dietz et al., 2025; Balog et al., 2025).
@@
This work treats LLMs not just as tools for text classification, but as participants that can imitate the subjective and sometimes inconsistent judgments made by humans. We introduce a practical method based Krippendorff’s $\alpha$, bootstrapping and TOST to test whether an LLM can blend into a group of human annotators without being identified. This approach requires only a small number of annotation items and functions as a domain-specific adaptation of the Turing Test. It supports early-stage evaluation on a small sample to determine the suitability of LLMs for large-scale annotation. We apply it to a real-world classification task and examine the results. Our key contributions are:
@@
  • We demonstrate the application of our methodology on two datasets—MovieLens 100K and PolitiFact—showing that the LLM is statistically indistinguishable from human annotators in MovieLens 100K ($p=0.004$) but not in PolitiFact ($p=0.155$), revealing important task-dependent differences.
@@ 2.1. Text Classification
Text classification is a core task in NLP. It helps organize unstructured text from sources like messages, documents, and websites. To make sense of all this text, researchers and developers use models that automatically classify text into categories such as topic, sentiment, or author identity (Siino et al., 2025). In addition, labeled data is essential for building and testing models, with human annotations serving as the “ground truth” (Nasution and Onan, 2024; Siino et al., 2025). Manual labeling is often costly, slow, and inconsistent, making automation preferable. Models are typically trained on a subset of labeled data and evaluated by comparing their predictions to human annotations.
Recently, LLMs have shown impressive performance across many NLP tasks  (Brown et al., 2020), raising the possibility of replacing traditional supervised models or even human annotators. However, this shift is far from complete (Movva et al., 2024). Many domain-specific real-world applications, such as medical coding, legal triage, or sentiment analysis, continue to rely on traditional classifiers trained on large datasets labeled by experts. This is because LLMs alone often lack the nuanced understanding required for these tasks. As a result, substantial human annotation remains essential for developing and validating reliable models. In areas where complex human judgment is critical, direct human involvement remains indispensable (Christoforou et al., 2025).
@@ 2.2. Human Annotation and Generative AI
Human annotation comes with challenges, most notably cost and consistency (Devillers et al., 2005). Text classification models rely on large amounts of labeled data, and hiring workers to do all the labeling is not always feasible. That is why crowdsourcing has become a solution. Platforms that connect requesters with crowd workers make it possible to outsource labeling tasks, such as relevance judgments, sentiment tagging, and topic categorization, to non-experts (Alonso and Mizzaro, 2012; Kutlu et al., 2018a). This approach has been crucial in building datasets needed to train machine learning models.
Crowdsourcing was once viewed as a flexible and empowering option for workers. However, it is often criticized as invisible, low-paid labor that supports modern AI behind the scenes. Given these concerns, researchers have started asking: Can generative AI (GenAI), especially LLMs, step in and take over some of these annotation tasks? Some early findings suggest that LLMs tend to perform well on straightforward tasks, such as summarization or basic sentiment analysis (Christoforou et al., 2025). But when the task requires more nuanced judgment—like interpreting sarcasm, ambiguity, or subtle context—human annotators still outperform the machines (Nasution and Onan, 2024).
So, the real question is not just whether GenAI can get the “right” answer. It is whether its decisions reflect the kinds of judgments humans would make, especially in cases where there is no single correct label. One study (Faggioli et al., 2023) examined the ability of LLMs to handle relevance judgments, a task where subjectivity plays a role. While LLMs showed some ability to mimic human responses, they were not consistent or nuanced enough to fully replace human workers.
@@ 2.3. Existing Evaluation Method
This raises an important question: if an LLM’s annotation looks like a human’s, does that mean it is truly the same, or are we missing differences in how judgments are made? In many tasks, there is no single “correct” answer; human judgments are often subjective, context-dependent, and inconsistent over time (Bailey et al., 2008; Faggioli et al., 2023). To better reflect this, we propose a new approach: instead of checking if an LLM agrees with the crowd, we ask whether it can blend into the crowd—becoming statistically indistinguishable from human annotators, and that we call it a “Statistical Turing Test”.
@@ 2.4. Human Judgment
This inconsistency can be attributed to the variability in cue weights applied across different tasks. Prior research has shown that judgment consistency tends to decrease as the number of cues increases (Einhorn, 1971). In contrast, LLMs often display greater consistency in annotation tasks (Faggioli et al., 2023), likely due to more stable internal representations of cues and weights. To evaluate whether the LLM’s cue integration falls within an acceptable threshold of variability, we use inter-annotator agreement (IAA) as a proxy for measuring consistency in annotation judgments.
@@ 2.5. Inter-Annotator Agreement (IAA)
Researchers who rely on hand-labeled data, where items are manually labeled with categories for empirical analysis or model development, must demonstrate that the labeling process is reliable (Artstein and Poesio, 2008). A fundamental assumption in annotation methodology is that the data are considered reliable when multiple annotators agree on the labels assigned, to a degree appropriate for the objectives of the study (Krippendorff, 2018; Craggs and Wood, 2004).
Consistent agreement among annotators suggests that they share a common understanding of the annotation guidelines, and thus can be expected to apply those guidelines consistently. IAA is a metric used to quantify this consistency. In multi-people annotation settings, annotators may have varied backgrounds and limited domain expertise. IAA helps determine whether labels are trustworthy and whether a task is clearly defined or inherently subjective. High agreement indicates clear instructions and straightforward data, while low agreement may reveal task ambiguity, multiple valid interpretations, or inconsistent annotator behavior.
@@ 3. Methodology
Building on the idea discussed in Section 2.5, consistent labeling by multiple human annotators suggests that they share a common understanding of the annotation guidelines and apply them reliably (Artstein and Poesio, 2008). Inspired by the logic of the Turing Test (Turing, 2009), we propose a methodology that uses inter-annotator agreement to evaluate whether an LLM can function as an individual annotator. That is, if it can serve as a substitute for a human in the annotation process.
@@ 3.1.1. Motivation: Substituting Human Annotators
Krippendorff’s $\alpha$ measures the extent of agreement among annotators using the formula:
$\alpha=1-\frac{D_{o}}{D_{e}}$
where $D_{o}$ is the observed disagreement and $D_{e}$ is the expected disagreement by chance.
@@ 3.1.3. Measuring Agreement Across Substitutions
We begin by computing Krippendorff’s $\alpha$ for the original group of $i$ human annotators. We denote this baseline agreement as:
$\alpha_{1}=\text{Krippendorff's $\alpha$ for human group}$
Next, we compute $\alpha$ for each of the $i$ modified groups where one human annotator is replaced by the LLM. These are denoted as:
$\alpha_{j}=\text{Krippendorff’s $\alpha$ for substitution group}j=2,\dots,i+1$
@@ 3.1.4. Paired Bootstrap for Variability Estimation
Let $\alpha_{1}^{(1)},\alpha_{1}^{(2)},\dots,\alpha_{1}^{(B)}$ represent the $\alpha$ values computed for the original human group across the $B$ bootstrap samples.
(1) $\alpha_{1}=\left\{\alpha_{1}^{(1)},\alpha_{1}^{(2)},\dots,\alpha_{1}^{(B)}\right\}$
For each LLM-substituted group $j=2,\dots,i+1$, we similarly compute a distribution of alpha values:
(2) $\alpha_{j}=\left\{\alpha_{j}^{(1)},\alpha_{j}^{(2)},\dots,\alpha_{j}^{(B)}\right\}$
@@ 3.2. Equivalence Testing with TOST
The null hypothesis is that the difference exceeds the equivalence margin:
$H_{0}:|x_{1}-x_{2}|>\Delta_{\text{equiv}}$
We reject $H_{0}$ if both $t_{1}$ and $t_{2}$ fall within the critical region for their respective one-sided tests, thereby concluding that the observed difference is within an acceptable range of equivalence.
@@ 3.2.1. Equivalence Margin Definition
The equivalence margin $\Delta_{\text{equiv}}$ is a threshold below which the difference is considered negligible. We estimate this margin empirically based on natural variability among human annotators:
$\Delta_{\text{equiv}}=(\alpha_{a}-\alpha_{b})\cdot\text{fraction}$
where:
@@ 3.3.2. Annotator Group Size
We analyze how Krippendorff’s $\alpha$ changes when one human annotator is replaced by an LLM. Let $i$ be the number of annotators, $n$ the number of items, and $\ell_{ak}$ the label from annotator $a$ on item $k$. As mentioned in Section 2.6, Krippendorff’s $\alpha$ is defined as:
$\alpha=1-\frac{D_{o}}{D_{e}},$
where $D_{o}$ is the average observed pairwise disagreement and $D_{e}$ is the expected disagreement under random labeling based on marginal label distributions.
@@ 4.1. Dataset
MovieLens 100K (Harper and Konstan, 2015) Movie ratings 1,682
WebCrowd25k (Kutlu et al., 2018b) IR relevance $\approx$4,500
TREC-8 Re-assessments (Roitero et al., 2021) IR relevance 4,269
@@
MBIC (Spinde et al., 2021) Media bias $\approx$2,600
CoQA (Reddy et al., 2019) Conversational QA 8,000
POPQUORN (Pei and Jurgens, 2023) Offensive–QA $\approx$5,500
D3CODE (Mostafazadeh Davani et al., 2024) Cross-cultural offensiveness detection $\approx$4,500
@@
CODA-19 (Huang et al., 2020) COVID-19 abstract section labeling 10,966
We compiled a list of publicly available datasets that provide multiple annotations per item along with identifiable annotator IDs, enabling us to trace which individual labeled each item and to perform annotator-level substitutions. These datasets are summarized in Table 1. To fit our evaluation experiments, we filtered the datasets using a set of selection criteria defined by our evaluation methodology. The code for this filtering process is available. [^2] Our filtering code can be found in https://github.com/peanutH/LLM-evaluation Specifically, we required that the dataset contain two disjoint groups of annotators, each of size $i$, such that there are at least 80 items, with each item annotated by at least two annotators from each group. After data filtering, two datasets satisfy our criteria: MovieLens 100K (Harper and Konstan, 2015) and PolitiFact (Roitero et al., 2020).
@@
The original PolitiFact dataset collected by Wang (2017) contains 12,000 statements produced by U.S. politicians, each statement is labeled by an expert judge on a six-level scale for the statement’s truthfulness. Roitero et al. (2020) selected 120 statements, 20 for each truth level, related to COVID-19 from the original PolitiFact dataset. Then, workers were recruited from Amazon Mechanical Turk to annotate each statement. Overall, each statement was annotated by 10 workers over three different scales: three-level, six-level, one-hundred level. In this work, we use the three-level dataset.
To obtain the LLM annotations, we used GPT-4o mini to annotate the same set of items selected for this experiment, as described in Section 4.2. We designed the prompt, see Figure 4, to follow the same annotation guidelines provided to human annotators.
–SYSTEM–
You are an average movie watcher. Rate each movie from 1 to 5 based on how much you liked it overall. Consider the story, acting, and overall enjoyment.
–USER–
Respond with a list of ratings, one for each movie, in the same order as presented. Only include the numeric ratings and nothing else.
@@
¡LIST_OF_MOVIES¿
Figure 4. LLM prompt for generating annotations for the MovieLens dataset. Further information is available on the paper’s GitHub.